### v1

# Efficiency-Effectiveness Reranking FLOPs for LLM-based Rerankers

Abstract
Large Language Models (LLMs) have recently been applied to reranking tasks in information retrieval, achieving strong performance. However, their high computational demands often hinder practical deployment. Existing studies evaluate the efficiency of LLM-based rerankers using proxy metrics such as latency, the number of forward passes, input tokens, and output tokens. However, these metrics depend on hardware and running-time choices (e.g., parallel or not, batch size, etc), and often fail to account for model size, making it difficult to interpret and obscuring the evaluation of the efficiency-effectiveness tradeoff. To address this issue, we propose E2R-FLOPs, for LLM-based rerankers: ranking metrics per PetaFLOP (RPP) for relevance per compute and queries per PetaFLOP (QPP) for hardware-agnostic throughput. Companied with the new metrics, an interpretable FLOPs estimator is built to estimate the FLOPs of an LLM-based reranker even without running any experiments. Based on the proposed metrics, we conduct comprehensive experiments to evaluate a wide range of LLM-based rerankers with different architecture, studying the efficiency-effectiveness trade-off and bringing this issue to the attention of the research community.
Efficiency-Effectiveness Reranking FLOPs for LLM-based Rerankers
Zhiyuan Peng Santa Clara University Santa Clara, CA zpeng@scu.edu                        Ting-ruen Wei Santa Clara University Santa Clara, CA twei2@scu.edu                        Tingyu Song Independent Researcher Beijing, China songtingyu220@gmail.com
Yilun Zhao Yale University New Haven, CT yilun.zhao@yale.edu                        Yi Fang Santa Clara University Santa Clara, CA yfang@scu.edu
1 Introduction
A typical search system balances efficiency and quality with a two-stage pipeline: a lightweight retriever retrieves hundreds of documents from a vast corpus, prioritizing efficiency, and then a more powerful but computationally expensive reranker refines their order. Thanks to the rapid progress of LLMs Brown et al. (2020); Grattafiori et al. (2024); Team et al. (2023), LLM-based rerankers have achieved impressive gains in reranking metrics, such as NDCG; however, these gains often come at the cost of substantial computational expense, making them difficult to deploy at scale in production. This underscores the need for evaluation metrics that consider not only reranking quality but also computational efficiency.
Existing studies evaluate the efficiency of LLM-based rerankers using proxies such as latency Jin et al. (2025), the number of LLM calls (i.e., forward passes) Zhuang et al. (2024), and input and output token usage Chen et al. (2025). However, these metrics lack computational granularity needed to distinguish differences in internal compute per token or per model call. Specifically, latency is heavily dependent on hardware and runtime choices (GPU vs. CPU, batch size, parallelism), making it an inconsistent basis for comparing algorithms across studies. The number of LLM calls ignores the model size: a single call to a 70B LLM costs orders of magnitude more compute than a call to a 3B model, yet both appear identical under this metric. Similarly, token usage overlooks the model size and is difficult to interpret as the cost of the input token and the output token can be different.
Inspired by the scaling law in LLMs that studies the connection between total compute and performance  Kaplan et al. (2020), we employ floating-point operations (FLOPs) as a fundamental measure of cost for each forward pass or LLM call. The total number of FLOPs required by a model to rerank documents is a hardware-agnostic, intrinsic metric of computational work  Sukthanker et al. (2024). Building on this insight, we introduce E2R-FLOPs, Efficiency- Effectiveness Reranking FLOPs for LLM-based rerankers: ranking metrics per PetaFLOP (RPP) for relevance per compute and queries per PetaFLOP (QPP) for hardware-agnostic throughput. The proposed metrics thus enable fair comparisons between methods that might utilize different LLMs, reranking algorithms, and running-time choices. Companied with the proposed metrics, an interpretable FLOPs estimator was built to estimate the FLOPs of an LLM-based reranker even without running any experiments.
Based on the proposed metrics, we conduct comprehensive experiments to evaluate a wide range of LLM-based rerankers, examining the efficiency-effectiveness trade-off and drawing attention to this issue within the research community. Our key contributions include:
  • We derive a closed-form, interpretable formula for the FLOPs of LLM-based rerankers and provide an open-source calculator covering up-to-date models and decoding settings.
  • We propose two efficiency-effectiveness metrics: RPP for relevance per compute and QPP for hardware-agnostic throughput.
  • We conduct the first large-scale study of the efficiency–effectiveness trade-off in LLM-based rerankers, bring this issue to the attention of the research community.
  • All code, data, and the FLOPs estimator are publicly released for reproducible research on computationally efficient rerankingEfficiency-Effectiveness Reranking FLOPs  [^1] https://github.com/zhiyuanpeng/EER-FLOPs..
2 Related Work
2.1 LLM-based Rerankers
Based on how the documents are compared with each other, LLM-based rerankers can be categorized as pointwise, pairwise and listwise. Pointwise methods primarily compute the query-document relevance score by either the likelihood of generating the query conditioned on the document Ponte and Croft (2017); Zhuang and Zuccon (2021); Zhuang et al. (2021); Peng et al. (2024) or the normalized possibility of generating the “Yes” when prompting the LLM whether the query-document pair is relevant or not Liang et al. (2023); Nogueira et al. (2020). The ranking can be easily accomplished by sorting the relevance score of each document. Pairwise methods compare the relevance of a pair of documents to a given query and output the document ID of the more relevant one. To rank a list of documents, sorting Qin et al. (2024) and sampling Gienapp et al. (2022); Mikhailiuk et al. (2020) methods are proposed. Sorting uses the pairwise comparison to replace the comparison operation in sorting algorithms like bubble sorting and heap sorting, while the sampling method. Soring method is more efficient to get the top-K document as it does not need to compare all the pairs. Setwise Zhuang et al. (2024) extends pairwise by comparing more than two documents in one LLM call. Listwise methods directly output a ranked list of document IDs. Most of the existing listwise methods are zero-shot Ma et al. (2023) or few-shot Sun et al. (2023); Ma et al. (2023); Pradeep et al. (2023) prompting methods. Recently, researchers Zhang et al. (2025) have resorted to adopting reinforcement learning to fine-tune LLMs to generate reasoning followed by ranked document IDs to tackle reasoning-intensive tasks like BRIGHT Su et al. (2025). To get a full rank list, listwise methods usually adopt strategies like sliding-window Sun et al. (2023) and tournament ranking  Chen et al. (2025) to get a full rank list with a limited window size.
2.2 FLOPs Calculation
Several FLOPs profilers exist for deep learning models, but most are limited to standard forward passes and do not support token-level generation with KV-cache, which is essential for accurate LLM inference estimation. PyPAPI LOISON (2023) measures CPU-level FLOPs for general Python code but is not designed for PyTorch or GPU workloads. ptflops Sovrasov (2018) and fvcore facebookresearch (2019) compute FLOPs by running a model’s “forward” function, but do not support autoregressive decoding. DeepSpeed’s FLOPs profiler Rasley et al. (2020) and calflops MrYxJ (2023) both support the FLOPs of the decoding process, but they also require full forward execution. All existing tools require model execution and lack closed-form support for generation-aware FLOPs estimation. For reranking, prior studies utilize coarse FLOP estimates, e.g., double the total parameter count  Shao et al. (2025) or open-source tooling  Abdallah et al. (2025), which lack interpretability regarding the specific facts that impact the FLOPs count. Our work differs from theirs in that our FLOPs estimator is well-interpretable, and we propose new metrics to comprehensively evaluate the efficiency and effectiveness of LLM-based rerankers.
3 Method
Multi-head Attention Grouped-query Attention
Ops Parameters FLOPs per Token Parameters FLOPs per Token
Atten: QKV $n_{\text{layer}}d_{\text{model}}3d_{\text{attn}}$ $2n_{\text{layer}}d_{\text{model}}3d_{\text{attn}}$ $n_{\text{layer}}d_{\text{model}}\,(1+2n_{\text{kv}}/n_{q})\,d_{\text{attn}}$ $2n_{\text{layer}}d_{\text{model}}\,(1+2n_{\text{KV}}/n_{Q})\,d_{\text{attn}}$
Atten: O $n_{\text{layer}}d_{\text{attn}}d_{\text{model}}$ $2n_{\text{layer}}d_{\text{attn}}d_{\text{model}}$ $n_{\text{layer}}d_{\text{attn}}d_{\text{model}}$ $2n_{\text{layer}}d_{\text{attn}}d_{\text{model}}$
Atten: Mask - $4n_{\text{layer}}d_{\text{attn}}d_{\text{model}}$ $4n_{\text{layer}}n_{\text{ctx}}(n_{\text{KV}}/n_{Q})d_{\text{attn}}$
Feedforward $n_{\text{layer}}2d_{\text{model}}d_{\text{dff}}$ $2n_{\text{layer}}2d_{\text{model}}d_{\text{dff}}$ $n_{\text{layer}}\,2d_{\text{model}}d_{\text{dff}}$ $2n_{\text{layer}}\,2d_{\text{model}}d_{\text{dff}}$
Table 1: FLOP count for the attention mechanism for multi-head attention and grouped-query attention
3.1 FLOPs Estimator
We parameterize a Transformer with four hyper-parameters: the number of layers $n_{\text{layer}}$, the residual-stream width $d_{\text{model}}$, the hidden size of the feed-forward block $d_{\text{ff}}$, and the dimension of attention output $d_{\text{attn}}$ which is the dimension of Q, K, V projections before splitting into multiple heads and by default $d_{\text{attn}}=d_{\text{model}}$. Because decoder-only and encoder–decoder designs dominate LLM-based rerankers, we derive estimates for both. To keep the analysis general yet concrete, we adopt the baseline decoder-only configuration of Kaplan et al. (2020) and the T5 encoder–decoder architecture (Raffel et al., 2020). In a typical reranking call, the model receives a prompt (the context, denoted ctx) and produces an output sequence (opt). The prompt concatenates a task-specific prefix $p$, the query $q$, and a list of $w$ documents, giving a length $n_{\text{ctx}}$. The generated sequence has length $n_{\text{opt}}$.
3.1.1 Decoder-only
Following Kaplan et al. (2020), we ignore sub-leading costs such as nonlinearities, biases, and layer-normalization. The number of attention and feedforward relevant parameters $N_{\text{dec}}$ is:
$\begin{split}N_{\text{dec}}&\approx 2d_{\text{model }}n_{\text{layer }}\left(2d_{\text{attn }}+d_{\text{ff}}\right)\end{split}$ (1)
Given an LLM with KV cache enabled, we now compute the FLOPs of generating a sequence named opt (short for “output”), consisting of $n_{\text{opt}}$ tokens, conditioned on a prompt ctx (short for “context”) of length $n_{\text{ctx}}$. The context includes a task-specific prompt $p$, a query $q$, and a list of $w$ documents. Each token within ctx requires $2N_{\text{dec}}+4n_{\text{layer }}n_{\text{ctx}}d_{\text{attn}}$ FLOPs, where $2N_{\text{dec}}$ comes from the fact that each parameter in $N_{\text{dec}}$ has one addition and one multiplication operation and $4n_{\text{layer }}n_{\text{ctx}}d_{\text{attn}}$ is taken by the basic multi-head attention operation  Kaplan et al. (2020). The total FLOPs for $n_{\text{ctx}}$ tokens $C({\text{ctx}})$ is:
$C\left(\text{ctx}\right)=2N_{\text{dec}}n_{\text{ctx}}+4n_{\text{layer }}n_{\text{ctx}}^{2}d_{\text{attn}}$ (2)
When generating token $\text{opt}_{i}$, the total sequence length seen by LLM is $n_{ctx}+(i-1)$ and the FLOPs for token $i$ is:
$C\left(\text{opt}_{i}\right)=2N_{\text{dec}}+4n_{\text{layer }}\left[n_{\text{ctx}}+(i-1)\right]d_{\text{attn}}$ (3)
The FLOPs of generating $n_{\text{opt}}$ tokens is computed by summing over all the $n_{\text{opt}}$ tokens:
$\begin{split}C({\text{opt}})&=2N_{\text{dec}}n_{\text{opt}}\\ &+2n_{\text{layer }}d_{\text{attn}}\left[2n_{\text{opt}}n_{\text{ctx}}+n_{\text{opt}}(n_{\text{opt}}-1)\right]\end{split}$ (4)
The total FLOPs of taking prompt ctx and generating opt is:
$C(\text{ctx}+\text{opt})=C(\text{ctx})+C(\text{opt})$ (5)
For LLM-based reranker, $n_{\text{ctx}}$ consists of task-specific prompt $p$, query $q$, and a list of $w$ documents. By approximating the length of each document as the average document length $L_{\text{doc}}$, $n_{\text{ctx}}$ can be estimated as:
$n_{\text{ctx}}=n_{p}+n_{q}+wl_{\text{doc}}$ (6)
Suppose $n_{\text{Q}}$ represents the number of heads for Q and $n_{\text{KV}}$ denotes the number of heads for K and V. Compare to multi-head attention, the number of parameters and the FLOPs per token changed accordingly as shown in Table 1. The equations are rewritten as:
$\begin{split}N_{\text{dec}}&\approx 2d_{\text{model }}n_{\text{layer }}\left((1+\frac{n_{\text{KV}}}{n_{\text{Q}}})d_{\text{attn }}+d_{\text{ff}}\right)\end{split}$ (7)
$C\left(\text{ctx}\right)=2N_{\text{dec}}n_{\text{ctx}}+4n_{\text{layer }}n_{\text{ctx}}^{2}\frac{n_{\text{KV}}}{n_{\text{Q}}}d_{\text{attn}}$ (8)
$\begin{split}C({\text{opt}})=\ &2N_{\text{dec}}n_{\text{opt}}+2n_{\text{layer}}\frac{n_{\text{KV}}}{n_{\text{Q}}}d_{\text{attn}}\cdot 2n_{\text{opt}}n_{\text{ctx}}\\ &+2n_{\text{layer}}\frac{n_{\text{KV}}}{n_{\text{Q}}}d_{\text{attn}}\cdot n_{\text{opt}}(n_{\text{opt}}-1)\end{split}$ (9)
3.1.2 Encoder-Decoder
Decoder-only LLMs, such as GPT Radford et al. (2019), do not include encoder-decoder attention and therefore share a similar structure with the encoder component of encoder-decoder models. The main difference lies in the attention masking strategy, which, however, does not affect the FLOPs required to process the prompt. Although decoder-only models are designed to compute attention only over previous tokens, in practice, they compute full self-attention (e.g., $Q_{\text{prompt}}\times K_{\text{prompt}}$) and apply a causal mask to prevent attending to future tokens. So, for encoder-decoder LLMs, the FLOPs of consuming prompt ctx is the same as that of encoder-only LLMs:
$C\left(\text{ctx}\right)=2N_{\text{enc}}n_{\text{ctx}}+4n_{\text{layer }}n_{\text{ctx}}^{2}d_{\text{attn}}$ (10)
Where $N_{\text{enc}}$ is same as $N_{\text{dec}}$ in Equation 1. The decoder has a different attention mechanism from that of the encoder, as it employs an encoder-decoder attention mechanism followed by self-attention. In an encoder–decoder model each decoder layer must, once per prompt, project the encoder outputs to cross-attention keys and values. This setup cost is calculated as:
$C_{\mathrm{cross\text{-}KV}}=4\,n_{\text{layer}}\,n_{\text{ctx}}\,d_{\text{model}}\,d_{\text{attn}}$ (11)
Even, it has two attentions, when generating token $\text{opt}_{i}$, it only goes through two Q projections, two O one K projection, and one V projections, as the left Q and K projections are for prompt “ntx”, so
$\begin{split}N_{\text{dec}}&\approx 2d_{\text{model }}n_{\text{layer }}\left(3d_{\text{attn }}+d_{\text{ff}}\right)\end{split}$ (12)
The total sequence length seen by self-attention is $(i-1)$ and the attention operation takes $4n_{\text{layer }}(i-1)d_{\text{attn}}$ FLOPs. The sequence length seen by the following encoder-decoder attention is fixed as $n_{\text{ctx}}$, which requires $4n_{\text{layer }}n_{\text{ctx}}d_{\text{attn}}$ for computing attention. The total attention FLOPs is $4n_{\text{layer }}(n_{\text{ctx}}+i-1)d_{\text{attn}}$ which is the same as that of decoder-only models as shown in the right part of equation 3 and thus the total FLOPs at generating token $\text{opt}_{i}$ is also same as equation 3 and the only difference is that the value of $N_{dec}$ is different. Similary, the $C({\text{opt}})$ is same as equation 4 but with a different value of $N_{\text{dec}}$. The cost of encoder-decoder model is:
$C(\text{ctx}+\text{opt})=C(\text{ctx})+C_{\mathrm{cross\text{-}KV}}+C(\text{opt})$ (13)
3.2 Metrics
We report two FLOPs-normalized metrics to compare different LLM-based rerankers, thereby effectively evaluating the effectiveness-efficiency tradeoff without being tied to a specific hardware.
3.2.1 Ranking metrics per PetaFLOP (RPP)
$\text{RPP}=\frac{m(q)}{C_{\text{q}}/10^{15}}$ (14)
where $m(q)$ can be any ranking metric for query $q$ (e.g., NDCG, MRR, MAP). RPP therefore expresses ranking metrics per peta-FLOP, the larger the value, the higher the ranking quality is obtained for a fixed compute budget.
3.2.2 Queries per PetaFLOP (QPP)
$\text{QPP}=\frac{1}{AVG_{C_{\text{q}}}/10^{15}}$ (15)
QPP measures throughput: how many queries can be processed with one peta-FLOP. Together, RPP and QPP trace a method’s efficiency–effectiveness frontier: RPP weights quality per compute, while QPP captures raw FLOPs-normalized throughput.
4 Experiment Setup
Following the setwise setup in Zhuang et al. (2024), we utilize the Flan-T5 Chung et al. (2024) as the backbone for most of the LLM-based rerankers except for IRL and Tourrank as these two methods require a longer context than Flan-T5’s input limitation allows. For IRL and Tourrank, we employ the Llama-3.1-8B-Instruct model. To compare our estimated FLOPs with those reported by open-source packages, we also include Qwen2.5 Yang et al. (2025) (3B, 7B, 14B), which implements group query attention instead of multi-head attention. For all LLM-based rerankers, we report their performance using the new metrics on the TREC-DL19 and DL20 datasets Craswell et al. (2020). The top 100 documents are retrieved using Pyserini’s BM25 Lin et al. (2021).
For FLOPs measurement, we utilize DeepSpeed’s FLOPs profiler Rasley et al. (2020) and calflops MrYxJ (2023) profilers and find identical results from both sources.
TREC DL 2019 TREC DL 2020
Methods NDCG #LLM In Out #FLOPs RPP QPP NDCG #LLM In Out #FLOPs RPP QPP
BM25 .506 - - - - - - .480 - - - - - -
Flan-t5-large pointwise.qlm .557 100 152.12 0.0 0.009 61.89 111.1* .567 100 152.85 0.0 0.009 63.0 111.1*
pointwise.yes_no .654 100 161.12 0.0 0.009 72.67* 111.1* .615 100 161.85 0.0 0.009 68.33* 111.1*
listwise.generation .561 245 486.21 10.54 0.076 7.38 13.16 .547 245 488.28 10.04 0.076 7.2 13.16
listwise.likelihood .669 245 384.49 0.0 0.058 11.53 17.24 .626 245 388.61 0.0 0.058 10.79 17.24
pairwise.allpair .666 9900 304.48 5.0 1.865 0.36 0.536 .622 9900 304.47 5.0 1.865 0.33 0.536
pairwise.heapsort .657 230.3 455.72 10.0 0.066 9.95 15.15 .619 226.8 459.62 10.00 0.066 9.38 15.15
pairwise.bubblesort .636 844.2 451.77 10.0 0.242 2.63 4.132 .589 778.5 459.03 10.0 0.227 2.59 4.405
setwise.heapsort .670 125.4 322.65 5.0 0.025 26.8 40.0 .618 124.2 325.5 5.0 0.025 24.72 40.0
setwise.bubblesort .678 460.5 320.9 5.0 0.091 7.45 10.99 .624 457.4 325.63 5.0 0.092 6.78 10.87
Flan-t5-xl pointwise.qlm .542 100 152.12 0.0 0.034 15.94 29.41 .542 100 152.85 0.0 0.034 15.94 29.41
pointwise.yes_no .650 100 161.12 0.0 0.036 18.06 27.78 .636 100 161.85 0.0 0.036 17.67 27.78
listwise.generation .569 245 486.38 11.87 0.282 2.02 3.546 .547 245 489.04 11.49 0.283 1.93 3.534
listwise.likelihood .689 245 385.49 0.0 0.216 3.19 4.629 .672 245 388.97 0.0 0.218 3.08 4.587
pairwise.allpair .713 9900 298.33 5.0 6.826 0.1 0.146 .682 9900 297.93 5.0 6.817 0.1 0.147
pairwise.heapsort .705 241.9 455.26 10.0 0.259 2.72 3.861 .692 244.3 455.76 10.0 0.262 2.64 3.817
pairwise.bubblesort .683 886.9 451.42 10.0 0.942 0.73 1.061 .662 863.9 457.18 10.0 0.930 0.71 1.075
setwise.heapsort .693 129.5 321.74 5.0 0.096 7.22 10.42 .678 127.8 325.27 5.0 0.096 7.06 10.42
setwise.bubblesort .705 446.9 335.53 5.0 0.346 2.04 2.890 .676 463.5 326.32 5.0 0.349 1.94 2.865
Flan-t5-xxl pointwise.qlm .506 100 152.12 0.0 0.135 3.75 7.407 .492 100 152.85 0.0 0.136 3.62 7.352
pointwise.yes_no .644 100 161.12 0.0 0.143 4.5 6.993 .632 100 161.85 0.0 0.144 4.39 6.944
listwise.generation .662 245 487.08 11.53 1.105 0.6 0.904 .637 245 489.60 11.05 1.110 0.57 0.901
listwise.likelihood .701 245 385.87 0.0 0.851 0.82 1.175 .690 245 389.73 0.0 0.860 0.8 1.162
pairwise.allpair .699 9900 282.32 5.0 25.51 0.03 0.039 .688 9900 282.32 5.0 25.51 0.03 0.039
pairwise.heapsort .708 239.4 456.98 10.0 1.010 0.7 0.990 .699 240.5 458.26 10.0 1.017 0.69 0.983
pairwise.bubblesort .679 870.5 453.06 10.0 3.642 0.19 0.275 .681 842.9 459.56 10.0 3.577 0.19 0.279
setwise.heapsort .706 130.1 323.43 5.0 0.383 1.84 2.610 .688 128.1 325.01 5.0 0.379 1.82 2.638
setwise.bubblesort .711 468.3 321.94 5.0 1.375 0.52 0.727 .686 467.9 326.37 5.0 1.393 0.49 0.717
L3.1 IRL .649 2 4469.12 0.0 0.096 6.76 10.42 .639 2 4556.31 0.0 0.098 6.52 10.20
Tourrank .757 130 1651.62 27.91 2.274 0.33 0.440 .777 130 1659.93 26.79 2.284 0.34 0.438
Table 2: Results on TREC DL. All the methods re-rank BM25 top 100 documents. #LLM represents average number of LLM call per query for reranking 100 documents. “In” and “Out” denote the average input tokens and output tokens per LLM call. #FLOPs is the estimated FLOPs per query for reranking 100 documents. Bold value is the best within each LLM and stared value is the best across different LLMs. “L3.1” represent Llama-3.1-8B-Instruct model. We report NDCG@10 for the NDCG metric.
5 Experimental Results and Analysis
We conduct extensive experiments to study four key research questions. Our results and analysis are as follows:
Q1: What are the LLM-based reranker performances under the new metrics?
Table 2 reports the efficiency–effectiveness trade-offs of a broad set of LLM-based rerankers under the proposed RPP and QPP metrics. Overall, these systems perform poorly once computation is taken into account. Across all LLMs and methods, TourRank with Llama-3.1-8B-Instruct secures the highest NDCG on both DL19 and DL20 at the cost of almost the lowest RPP and QPP. Beyond this result, several additional insights emerge:
Pointwise methods dominate the RPP and QPP metrics across different LLMs and datasets. pointwise.yes_no of Flan-T5-large yields the highest RPP of 72.67 (DL19) and 68.3 (DL20) and achieves the maximum QPP of 111 queries/PetoFLOPs. The variant pointwise.qlm obtains the similar QPP metric but 5 10 worse RPP points. These methods obtain 10% to 30% relative NDCG gains over the baseline BM25 with negligible FLOPs consumption compared with other LLM-based rerankers.
Scaling up hurts efficiency far more than it helps effectiveness. Most LLM-based rerankers gain NDCG when moving from Flan-T5-large to Flan-T5-xl, yet see only marginal improvement from xl to xxl. For example, setwise. Heapsort rises from 0.670 to 0.693 and then to 0.706 in NDCG. Meanwhile, efficiency collapses: RPP plunges from 26.8 to 7.22 and then to 1.84, while QPP falls from 40.0 to 10.42 and finally to 2.61. In short, scaling boosts quality slowly but sacrifices RPP and QPP on a much larger scale.
Pairwise and listwise methods are intensely FLOP-hungry. Allpair sorting, though it delivers the highest NDCG on Flan-T5-xl (0.713), issues 9,900 LLM calls per query; its RPP collapses to around 0.10, and it processes barely 0.15 queries per peta FLOPs, making large-scale deployment impractical. Heapsort and bubblesort based variants cut the number of calls by roughly 90%, yet remain about orders of magnitude less efficient than pointwise methods on both RPP and QPP.
Q2: Do the estimated FLOPs reflect the measured FLOPs?
Figure 1 shows the relationship between the estimated and real FLOPs on DL19 for models of various sizes. The comparison contains both decoder-only and encoder-decoder architectures, providing a comprehensive view of scaling trends. In both cases, the estimated and real FLOP counts scale with the model size, reflecting the expected rise in computational cost with increasing model parameters. The linear pattern across models illustrates that the estimated FLOPs correlate linearly with the real FLOPs and is consistent across model families and architectural types, suggesting that our FLOPs estimator is accurate and reliable. The close alignment between the two quantities provides empirical justification for the FLOPs estimator described in Section 3, affirming its reliability as a proxy when real measurements are unavailable.
Q3: How does latency relate to the FLOP counts?
Figure 2 shows the relationship between latency and FLOP counts for two representative models: Qwen-7B (a decoder-only architecture) and Flan-T5-XXL (an encoder-decoder architecture), evaluated on the DL19 dataset. For both models, we observe that latency increases in accordance with the number of FLOPs. Importantly, the estimated FLOP counts exhibit a correlation with latency that closely mirrors the relationship between real FLOP counts and latency. This alignment indicates that our FLOPs estimator approximates computational cost accurately and can serve as a reliable predictor of real-world latency trends. This means that the estimator can be used to anticipate inference time without requiring direct hardware profiling, which is particularly useful when comparing models in a platform-agnostic setting or during early-stage architecture design.
Q4: What is the relationship between prompt length and FLOPs?
Figure 3 shows the relationship between prompt length and FLOPs. As expected, both the estimated and actual FLOPs increase with longer prompts, reflecting the greater computational cost required to process more input tokens. Notably, the estimated FLOPs show a strong correlation with prompt length that closely mirrors the pattern observed in the actual FLOPs. As the prompt becomes longer, the estimator reliably tracks the resulting increase in computation, consistent with what is observed empirically. This result provides additional validation for the robustness of our FLOPs estimator, demonstrating its ability to respond to changes in input length in the same way real FLOPs do.
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
Figure 1: Linear trends between estimated and real FLOPs for decoder (left) and encoder-decoder (right) models of various sizes on DL19. The same is observed for the DL20 dataset.
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
Figure 2: Latency in milliseconds increases with FLOPs on Qwen-7B (left) and Flan-T5-XXL (right)
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
Figure 3: FLOPs increases with prompt length for Qwen-7B (left) and Flan-T5-XL (right) on the DL19 dataset.
6 Conclusion and Future Works
Due to limitation of existing metrics in evaluating the efficiency-effectiveness tradeoff of large language models as rerankers, we propose two metrics RPP and QPP to quantify the model performance. In addition, we provide a calculator through a closed-form and interpretable formula to compute the FLOPs and validate this estimation through experiments on existing decoder-only and encoder-decoder model architectures. The estimated FLOP count exhibits a strong linear correlation with the actual measured values, allowing it to approximate real-world computational cost even without model execution. Future work includes conducting a linear regression between the real and estimated FLOP counts to refine our estimation and adapting to more advanced architectures.
Limitations
While our proposed FLOP-normalized metrics and estimator provide a hardware-agnostic and interpretable approach to evaluating the efficiency–effectiveness tradeoff of LLM-based rerankers, there are some limitations. The FLOP estimation relies on model architecture specifications and assumes consistent implementation across different frameworks, which may not hold in practice due to library-level optimizations or kernel differences. Although the estimator shows strong linear correlation with actual FLOP measurements, the approximation may be less accurate for models with more advanced architectures such as mixture-of-experts. While FLOPs offer a more stable proxy than latency or token counts, they do not capture other real-world constraints such as memory bandwidth, energy consumption, or inference-time variability under dynamic system loads.
Appendix A Appendix
A.1 BM25
$\sum_{i=1}^{n}\text{IDF}(q_{i})\frac{f(q_{i},D)(k+1)}{f(q_{i},D)+k\left(1-b+b\frac{|D|}{\text{avgdl}}\right)}$ (16)
BM25 Robertson et al. (2009) computes a relevance score between a query and a document with Equation 16 where IDF denotes the inverse document frequency, $f(q_{i},D)$ represents the frequency of the $i^{th}$ query token in the document, $|D|$ is the length of the document, avgdl is the average document length in the corpus, and $k$ and $b$ are hyperparameters. Assuming that the term frequencies and the inverse document frequencies are precomputed and do not contribute to the runtime FLOP count, the upper bound on the FLOPs required for BM25 scoring can be estimated as:
$C(\text{BM25})=11\cdot L_{Q}\cdot N_{D}$ (17)
where $L_{Q}$ is the length of the query and $N_{D}$ is the number of documents, which is 100 for reranking top-100 documents in our experiments. This represents an upper bound because it assumes that every query token appears in every document. In cases where a query token does not appear in a document, the corresponding numerator becomes zero, resulting in zero FLOPs for that term-document pair instead of 11.